OpenStudy (anonymous):

Which circles have an area of 36pie ft2? Choose all answers that are correct. A. a circle with a radius of 6 ft B. a circle with a diameter of 6 ft C. a circle with a radius of 12 ft D. a circle with a diameter of 12 ft

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Plug it in the formula: \(C = 2 \pi r \) \(36 \pi = 2 \pi r\) Divide pi to both sides: \(36 = 2r\) Divide 2 to both sides, what's 36 / 2? @lordoftanks

OpenStudy (anonymous):

the %100 correct answer would actualy be 3.385 for the radius...

OpenStudy (anonymous):

Can you divide 36 / 2? @lordoftanks

OpenStudy (triciaal):

area of 36pie ft2 formula for the area A = pi* r^2 @iGreen. gave the circumference C = 2*pi*r

OpenStudy (triciaal):